OpenStudy (anonymous):

ind the maximum and minimum values of the function f(x,y,z) = x^2 y^2 z^2 subject to the constraint x^2 + y^2 + z^2 = 400. I am so confused because I got complicated results. Please help!

ganeshie8 (ganeshie8):

did u use lagrange ?

ganeshie8 (ganeshie8):

Clearly the minimum value is 0, occurs when any one of the variable is 0

ganeshie8 (ganeshie8):

you can find the maximum value by using symmetry : max value occurs when \(x = y = z\) \(\large \implies 3x^2 = 400 \) \(\large \implies x = \pm \frac{20}{\sqrt{3}}\)

OpenStudy (anonymous):

I did it but homework page didn'T accepted it. I said (20/sqrt3,20/sqrt3,20/sqrt3) is maximum points. Website said " This is not a number" .

ganeshie8 (ganeshie8):

maximum value : 64000000/27 minimum value : 0

ganeshie8 (ganeshie8):

try those^